Circular No. 8704

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 73P/SCHWASSMANN-WACHMANN G. A. Blake and C. Salyk, California Institute of Technology; B. P.

Bonev, G. L. Villanueva, M. A. DiSanti, and M. J. Mumma, Goddard Space Flight Center; K. Magee-Sauer, Rowan University; and E. L. Gibb, University Missouri at St. Louis, report that component ‘C’ was observed on Apr. 7 and 16 UT using NIRSPEC at the W. M. Keck Observatory. On Apr. 7, H_{2}O (twelve lines, 2.9-μm hot-bands), C_{2}H_{6} (^{R}Q_{0} and ^{P} Q_{1} near 3.35 μm, ν_{7} band), and HCN (six lines near 3.0 μm, ν_{3} band) were detected. Sensitive upper limits for CH_{4} (R0 and R1 near 3.3 μm, ν_{3} band), CH_{3}OH (3.52 μm Q-branch, ν_{2} band), and H_{2}CO (3.59 μm Q-branch, ν_{1} band) were also obtained. H_{2}O (eight lines) yielded a rotational temperature of 55 ± 6 K; for HCN (six lines) the value was 52 K, and a rotational temperature of 55 K was assumed for other species. Production rates (10^{25} molecules/s) or upper limits (3σ) for Apr. 7 are as follows: H_{2}O (590 ± 60), C_{2}H_{6} (0.9 ± 0.2), HCN (1 ± 0.1), CH_{4} (< 1.5), CH_{3}OH (< 3), H_{2}CO (< 3). The preliminary abundance ratios are then H_{2}O : C_{2}H_{6} : HCN : CH_{4} : CH_{3}OH : H_{2}CO = 100 : 0.15 : 0.2 : < 0.25 : < 0.5 : < 0.5. On Apr. 16, the abundance ratios were: H_{2}O : HCN : CH_{4} : H_{2}CO = 100 : 0.2 : < 0.25 : 0.5 ± 0.13, respectively. The abundance ratios (relative to H_{2}O) of C_{2}H_{6}, HCN, CH_{4}, and CH_{3}OH in component ‘C’ of 73P are similar to those seen in C/1999 S4 (LINEAR), and are much lower than those of most other comets.

Further to IAUC 8703, three additional components to 73P have been observed on three or more nights and thus given designations (‘AN’, ‘AO’, and ‘AP’) and first announced on MPEC 2006-H26.

COMET P/1999 RE_{70} = (118401)

H. H. Hsieh and D. Jewitt, University of Hawaii (UH), report that two 120-s R-band images of this object (with a = 3.20 AU, e = 0.19), obtained on 2005 Nov. 26, show a tail < 7′′ long in p.a. 90^{o}; the images were made at the Gemini North Telescope during their survey of Themis-family objects under the suspicion that (118401) might show cometary activity along with other objects having similar main-asteroid-belt orbital elements [133P/Elst-Pizarro = (7968) and P/2005 U1]. Confirming observations from 2005 Dec. 22, 24, 25, 26, and 27 with the UH 2.2-m telescope also show the cometary appearance of (118401) — the Dec. 22 observation being made by J. Pittichová, and the rest by Hsieh; a co-added 9900-s image from Dec. 27 shows a faint coma with a fan-shaped tail ∼ 9′′ long in p.a. 90^{o}.

2006 April 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green